# GenArtist: Multimodal LLM as an Agent for Unified Image Generation and Editing

### v1

@@ 2 Related Work
AI agent. Large language models (LLMs), like ChatGPT, Llama touvron2023llama ; touvron2023llama2 , have demonstrated impressive capability in natural language processing. The involvement of vision ability for multimodal large language models (MLLMS), like LLaVA liu2024visual , Claude, GPT-4 gpt42023 , further enables the models to process visual data. Recently, LLMs begin to be adopted as agents for executing complex tasks. These works yang2023auto ; shen2023hugginggpt ; liu2023internchat apply LLMs to learn to use tools for tasks like visual interaction, speech processing, software development qian2023communicative , gaming meta2022human , APP use yang2023appagent or math xin2023lego . Recently, the idea of AI agents has also begun to be applied to image generation related tasks. For example, lian2023llm ; feng2024layoutgpt design scene layout with LLMs, wu2023self utilizes LLMs to assist self-correcting, wang2024divide ; yang2024mastering target at MLLMs in complex text-to-image generation problems, and qin2024diffusiongpt leverages LLM for model selection in the text-to-image generation task.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of our GenArtist and its unified ability through extensive experiments in image generation and editing. For image generation, we mainly conduct quantitative comparisons on the recent T2I-CompBench benchmark huang2023t2i . It is mainly about image generation with complex text prompts, involving multiple objects together with their own attributes or relationships. For image editing, we mainly conduct comparisons on the MagicBrush benchmark zhang2024magicbrush , which involves multiple types of text instructions, both single-turn and multi-turn dialogs for image editing. We choose GPT-4 gpt42023 as our MLLM agent. In quantitative comparative experiments, we constrain the editing tree to be a binary tree.

### v2

@@ 2 Related Work
AI agent. Large language models (LLMs), like ChatGPT, Llama touvron2023llama ; touvron2023llama2 , have demonstrated impressive capability in natural language processing. The involvement of vision ability for multimodal large language models (MLLMS), like LLaVA liu2024visual , Claude, GPT-4 gpt42023 , further enables the models to process visual data. Recently, LLMs begin to be adopted as agents for executing complex tasks. These works yang2023auto ; shen2023hugginggpt ; liu2023internchat apply LLMs to learn to use tools for tasks like visual interaction, speech processing, compositional visual tasks gupta2023visual , software development qian2023communicative , gaming meta2022human , APP use yang2023appagent or math xin2023lego . Recently, the idea of AI agents has also begun to be applied to image generation related tasks. For example, lian2023llm ; feng2024layoutgpt design scene layout with LLMs, wu2023self utilizes LLMs to assist self-correcting, wang2024divide ; yang2024mastering target at MLLMs in complex text-to-image generation problems, and qin2024diffusiongpt leverages LLM for model selection in the text-to-image generation task.
@@ 3.1 Planning Tree with Step-by-Step Verification
It is also worth mentioning that during verification, in addition to the accuracy of the generated images, the agent is also required to assess their aesthetic quality. If the overall quality is poor, the agent will utilize different generation tools or choose different random seeds to regenerate the images, in order to ensure their overall quality. Meanwhile, as an agent-centered system, the framework is also flexible in terms of human-computer interaction. During verification, human feedback can be appropriately integrated. By incorporating human evaluation and feedback on the overall quality of the images, the quality of the generated images can be further improved.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of our GenArtist and its unified ability through extensive experiments in image generation and editing. For image generation, we mainly conduct quantitative comparisons on the recent T2I-CompBench benchmark huang2023t2i . It is mainly about image generation with complex text prompts, involving multiple objects together with their own attributes or relationships. For image editing, we mainly conduct comparisons on the MagicBrush benchmark zhang2024magicbrush , which involves multiple types of text instructions, both single-turn and multi-turn dialogs for image editing. We choose GPT-4V gpt42023 as our MLLM agent. In quantitative comparative experiments, we constrain the editing tree to be a binary tree.
@@ 4.3 Ablation Study
Table 5: Ablation Study on the position-aware tool execution on T2I-CompBench.
[…]
w/o position information 0.4577 0.4083
w/ position information 0.5437 0.4499
Regarding position-aware tool execution, we list the corresponding ablation study in Tab. 5. We evaluate the performance on the spatial and complex aspects of T2I-CompBench, as these two aspects mainly involve position-sensitive text prompts for image generation. As multimodal large models are usually not sensitive to position information, the performance is limited without the inclusion of position information, only a slight improvement over the tool selection results. After introducing position information, which enhances spatial awareness, there is a significant improvement in both the spatial and complex aspects. This validates the reasonability of our design.
@@
4.4 Error Case Analysis
Figure 6: The error cases of GenArtist caused by the ability of editing tools (the left) or the wrong output of localization tools (the right).
We further analyze some error cases from our GenArtist in Fig. 6. As can be seen, sometimes, despite the agent correctly planning the specific execution of tools, the limitations of the tools themselves prevent correct execution, leading to incorrect results. For example, in the first case, it is required to add a very small blue cup. However, due to the lack of fine resolution ability in existing editing tools, the generated blue cup’s size is inaccurate. In addition, as shown in the second case, errors in the output of localization tools can also affect the final result. For instance, when asked to remove the lettuce in the middle of a sandwich, the segmentation model fails to accurately identify the part of the object, leading to the erroneous removal operation. Utilizing more powerful tools or incorporating some human feedback during the verification stage can effectively address this issue.
@@ 5 Conclusion
Acknowledgement
We gratefully acknowledge the support of Mindspore, CANN(Compute Architecture for Neural Networks) and Ascend AI Processor used for this research.